OpenStudy (anonymous):

Consider the function f(x) = x^2-25/x+4 IS the VA = -4 is the HA = +/- 5 is the OA = x

OpenStudy (anonymous):

x^2-x-12/x-4 Are these correct? disc = 4 x int = (-3,0) y int = (0,3) basic shape= straight line

OpenStudy (cruffo):

\[\frac{x^2-x-12}{x-4} = \frac{(x-4)(x+3)}{x-4} = x+3\] So, your stuff yep looks good.

OpenStudy (anonymous):

are both problems correct?

OpenStudy (cruffo):

Original post) Remember you will have either a HA or an OA but not both

OpenStudy (anonymous):

so there is no HA on that problem

OpenStudy (cruffo):

Good. No HA. VA looks good. To get the OA do long division. Looks like y = x + something

OpenStudy (anonymous):

| | x | + | 4 | | x | - | 4 x^2 | + | 0 | - | 25 x^2 | + | 4 x | | | | -4 x | - | 25 | | -4 x | - | 16 | | | | -9

OpenStudy (anonymous):

OA = x+4 or x-4 .. i forget

OpenStudy (cruffo):

Right, don't forget the "missing term" OA looks to be y = x-4

OpenStudy (anonymous):

thank you

OpenStudy (cruffo):

Cheers!